The bridal shop has net sales of $33,000 and returns of 1/12 of gross sales. If net sales equal gross sales minus returns, find gross sales.
Answer by midwood_trail(310)

The bridal shop has net sales of $33,000 and returns of 1/12 of gross sales. If net sales equal gross sales minus returns, find gross sales.
Why not set up an equation?
n = net sales
r = returns
g = gross
Here is your equation:
n = g - r
Let n = 33,000
Let r = 1/12
33,000 = g - (1/12)(g)
We can write (1/12)(g)as (g/12).
Here is your equation ready to be solved:
33,000 = g - (g/12)
Can you take it from here?
All you have to do is solve for g.
If not, write back.